SOLVE THE WORD PROBLEM. NEED TO USE 2 LET STATEMENTS. PLEASE SOLVE SOMPLETELY
A chemist has 6% and 15% solutions of sulfuric acid. How much of each solution should she mix to get 10 liters of a 9% solution?

Answer by Boreal(15235)

Let x= number of liters of 6% sulfuric acid solution.
Let y= number of liters of 15% sulfuric acid solution.
x+y=10 (units are liters)
.06*x + (.15)y=(.09)*10
y=10-x (substitution works well here). Elimination could be used.
.06x + (.15) (10-x)=0.9 (I multiplied out the right side)
.06x + 1.5 - 0.15x=0.9
Subtract 1.5 from both sides and combine x s You will get two negatives, one on each side of the equation, and they will cancel out, making the result positive.
-0.09x=-0.6
x=6.67 (repeating decimal) liters of 6%
y=3.33 liters of 15%
Check is done a different way:
9%, the desired amount, is 1/3 the way between 6 and 15%
Therefore, 2/3 of the solution must come from the 6% solution and 1/3 from the 15% solution